The gravitational attraction between two bodies varies inversely as the square of the distance between them. If the force of attraction is 64 pounds when the distance between the bodies is 9 feet, what is the distance between them when the force is 36 pounds?
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

F = k/d^2
64 = k/81
k = 81(64) = 5184
.......
F = 5184/d^2
F = 36lb
d = √(5184/36) = 12' is the distance between them

Answer by stanbon(75887)

The gravitational attraction between two bodies varies inversely as the square of the distance between them. If the force of attraction is 64 pounds when the distance between the bodies is 9 feet, what is the distance between them when the force is 36 pounds?
----
a = k/d^2
----
Solve for "k"::
64 = k/9^2
k = 9^2*64
----
Equation:
a = (9^2*64)/d^2
Your Problem::
36 = (9^2*64)/d^2
----
d^2 = (9^2*64)/(36)
d = (9*8)/6
distance = 12 ft
------------------
Cheers,
Stan H